# Multiple References with Meaningful Variations Improve Literary Machine Translation

### v1

@@ Abstract
While a source sentence can be translated in many ways, most machine translation (MT) models are trained with only a single reference. Previous work has shown that using synthetic paraphrases can improve MT. This paper investigates best practices for employing multiple references by analyzing the semantic similarity among different English translations of world literature in the Par3 dataset (Thai et al., 2022). We classify the semantic similarity between paraphrases into three groups: low, medium, and high, and fine-tune two different LLMs (mT5-large and LLaMA-2-7B) for downstream MT tasks. Across different models, holding the total training instances constant, single-reference but more source texts only marginally outperforms multiple-reference with half of the source texts. Moreover, using paraphrases of medium and high semantic similarity outperforms an unfiltered dataset (+BLEU 0.3-0.5, +COMET 0.2-0.9, +chrF++ 0.25-0.32). Our code is publicly available on GitHub. [^1] https://github.com/swsiwu/multi_ref_literary_MT
@@ 1 Introduction
In this paper, we investigate two questions: when holding the amount of fine-tuning data constant, is using multiple references better than just one target reference? What kind of paraphrase should be incorporated to achieve the best performance?
To answer these questions, we create datasets that vary in the number of source texts, the number of references per source text (single vs multiple references), and for multiple-reference datasets, vary in the level of semantic similarity among the references. We use Paragram-SP (Wieting et al., 2022), an efficient and effective model for measuring semantic similarity. We divide paraphrases into three groups based on their semantic similarity: low, medium, and high, where medium has the most interesting and meaningful variations among the references for a source text. We then fine-tune two LLMs, mT5-large and LLaMA-2-7B, using aforementioned datasets on MT tasks, and compare their performances with baseline and commercial models: No Language Left Behind (NLLB Team et al., 2022) and Google Translate.
@@
  • Using source texts with medium and high semantic similarity among references can outperform an unfiltered dataset even if the unfiltered dataset has more training instances. (Section 6.2)
  • On the other hand, training with medium and low semantic similarity among references will perform worse than an unfiltered dataset. (Section 6.2)
@@ 2 Related work
Khayrallah et al. (2020) use simulated paraphrases to improve MT performance on low-resource languages. They focus on the general effect of augmenting training data using synthetic paraphrases when training data is scarce. Our work further studies the semantic similarity of paraphrases on MT. It is worth noting that the languages present in our dataset are not low-resource languages, and instead of training a language model from scratch, we fine-tune two different types of LLMs: mT5 (an encoder-decoder model) and LLaMA-2-7B (an autoregressive LLM).
@@ 3 Measuring semantic similarity
Figure 2: The distribution of sim_p for the Par3 corpus’s source texts. Paraphrases with high scores are more similar semantically and syntactically. Upon manual inspection, we observe that the paraphrases in between the two blue vertical lines are the ones with meaningful variations: different enough to be interesting, but similar enough to still be a paraphrase.
@@
e.g. single vs double quotation marks, or word replacement Unfiltered
@@ 4 Datasets
We use #source to denote the number of unique source texts, and #instance to denote the total number of training instances, e.g. 3 source texts each of which has 3 reference translations, then #source = 3, and #instance = 9. #source and #instance for each dataset are in Table 2. Here we describe some representative datasets:
Medium All
#source = 49,650: This dataset uses multiple references for each source text. It contains all the paraphrases with sim_p $\in[0.45,0.85]$, of which we define to have the most meaningful variations (§3).
Medium
#source = 20,897: This dataset uses multiple references for each source text and has the same amount of training instances (#instance) as Single Ref #source = 49,650, but its #source is smaller. This dataset is created to find out the extent of the efficiency of training with Medium datasets. Similarly, we create Medium with #instance = 57,987 and 76,938 to compare with the corresponding single-reference datasets.
Unfiltered
#source = 49,650: This dataset uses multiple references for each source text, and it contains the same amount of source texts as Medium All, but its source texts are randomly selected regardless of their corresponding target texts’ sim_p.
Single Ref
#source = 49,650: This dataset has the same amount of source texts as Medium All, but each source text is only trained with one reference.
Medium+Low 10%:
This dataset uses multiple references for each source text, and it includes the entire Medium All dataset. Additionally, we randomly select 10% of the source texts from the lower sim_p range ($[0,0.45]$).
Medium+Low All:
Similar to Medium+Low 10%, but it includes all the source texts with sim_p in the lower range ($[0,0.45]$).
Medium+High 10%:
Similar to Medium+Low 10%, except the additional 10% source texts are from high sim_p range ($[0.85,1]$). The Medium+High datasets are to examine the impact of including less meaningful variations (high semantic similarity) when fine-tuning an MT model.
The source paragraphs and human references are from Par3 (Thai et al., 2022). [^2] Par3 dataset can be downloaded from https://github.com/katherinethai/par3 The source languages include Portuguese, Swedish, Italian, Czech, Japanese, Classical Chinese, Norwegian, German, Russian, and French. All references are only in English; In other words, the language pairs in this dataset are all to-English language pairs. We provide details about the composition of Medium all compared to Unfiltered All in Appendix 4.
@@ 5 Experiments
We are interested in the effect of varying the number of source text, the number of references, and the semantic similarity of reference translations on MT tasks in the literary domain. Using datasets described in Section 4, we fine-tune two different general-purpose LLMs for translation: mT5-large and LLaMA-2-7B.
@@ 5.1 Models
(Xue et al., 2021): an encoder-decoder model that covers 101 languages (Xue et al., 2021) including all the languages in Par3. It is not specifically trained on translation tasks but could be fine-tuned for a variety of downstream tasks including multilingual translation. It has about 1.2 billion parameters. We choose to fine-tune mT5-large for our translation tasks. [^3] https://huggingface.co/google/mt5-large
@@
(Touvron et al., 2023): an English-centric LLM with limited multilingual capabilities. Studies have shown that LLaMA-2 can translate high-resource non-English language with some fine-tuning due to the presence of the corresponding language in the training dataset. LLaMA-2 is also able to adapt to a new low-resource language with some fine-tuning. All the languages in this paper are in the LLaMA-2 training data for >= 0.05%, except that the Chinese in Par3 is classical Chinese, which shares mostly the same characters as modern Chinese but has different grammar. It would be interesting to observe the translation performance after fine-tuning on our datasets given that non-English data is only a minuscule part of LLaMA-2 training data.
@@ 5.2 Evaluation metrics
(Popović, 2017): An MT evaluation metric that is based on character and word-level n-gram recall and precision. We include this metric as Zouhar et al. (2024) find that COMET can struggle in domains outside WMT, e.g. literary domain.
#source #instance BLEU $\text{COMET}_{22}^{\text{DA}}$ chrF++
NLLB-1.3B 27.3 0.689 41.62
Google Translate 35.8 0.726 48.02
mT5 LLaMA-2 mT5 LLaMA-2 mT5 LLaMA-2
medium 20,897 49,650 29.8 30.0 71.3 72.8 43.20 44.06
single ref 49,650 49,650 29.9 30.1 71.5 70.3 43.44 42.08
unfiltered 49,650 116,689 31.3 30.8 71.8 72.7 44.17 44.52
medium all 49,650 118,003 31.5 30.3 71.9 71.6 43.81 43.76
medium 24,398 57,987 29.8 30.2 71.3 72.4 43.36 43.79
single ref 57,987 57,987 30.4 30.2 71.5 72.1 43.81 42.86
medium 32,357 76,938 30.6 28.2 71.6 70.2 43.73 41.12
single ref all 76,938 76,938 31.3 29.4 71.8 71.3 44.33 42.81
unfiltered 57,987 136,356 31.7 30.4 71.9 72.2 44.63 44.13
medium + low (10%) 57,987 137,811 31.1 29.4 71.7 71.6 44.16 41.29
medium + low (all) 59,072 140,407 31.1 30.1 71.7 71.7 44.11 43.64
unfiltered all 76,938 180,927 31.9 30.8 71.9 72.1 44.67 44.11
medium + high (10%) 57,987 136,870 32.0 29.2 72.0 70.3 44.77 42.39
medium + high (20%) 66,324 155,809 32.2 30.8 72.1 72.6 44.92 44.43
medium + high (all) 67,516 158,523 32.4 31.1 72.1 73.0 44.84 44.31
Table 2: Different datasets’ performance across different models. #source is the number of source texts, and #instance is the number of total training instances. For each MT metric, $\dagger$ denotes statistical significance for experiments using mT5, and $\ddagger$ denotes statistical significance for experiments using LLaMA-2-7B. We tested the statistical significance (p-value < 0.05) using the paired bootstrap resampling method (Koehn, 2004).
@@ 5.3 Training details
Given many LLMs already have strong language abilities and only require fine-tuning for MT downstream tasks, we consider fine-tuning as a more efficient approach than training from scratch, especially since our dataset is English-centric (all reference translations are in English) and these LLMs usually perform better when translating from non-English language into English.
For fine-tuning, each training instance is a source text and one of its human translations, and he training and validation datasets are randomized, e.g. (source text A, target human translation $A_{1}$), (source text B, target human translation $B_{3}$), (source text C, target human translation $C_{1}$), etc.
@@
For LLaMA-2-7B, we used LoRA-tuning (Hu et al., 2022) with Meta’s LLaMA-recipe. [^7] https://llama.meta.com/llama-downloads/, [^8] https://github.com/meta-llama/llama-recipes We structured each training instance as "source_text ###> target text" following the fine-tuning method from Zhang et al. (2023). We used a batch size of 16. Depending on the dataset size, training took 8 to 24 hours.
For both models, we use BLEU on the validation set for early stopping. We notice that unlike mT5, the LLaMA-2 fine-tuning hyperparameters impact the performance significantly (±$\pm 5$ BLEU score). We also discover that when the training data increases significantly, increasing global batch size seems to improve the performance. In this paper, we report the best performance we obtained with three different global batch sizes for LLaMA-2-7B.
@@ 6 Results and Discussion
Figure 3: Performance difference between different languages under different MT metrics when fine-tuning with mT5-large (1st row) and LLaMA-2-7B (2nd row). Darker dots are from Medium+High All, and lighter dots are from Unfiltered All. The bar chart show the amount of source-target-text pairs in Unfiltered All dataset.
@@ 6.1 Single reference vs multiple references
This is not, to our knowledge, an idea that has been previously explored. In Table 2, when n_instance is the same, Single Ref on average only marginally outperforms Medium across different models and metrics: ranging 0.0-0.7 in BLEU, (-)2.5-(+)1.1 in COMET, and (-)1.98-(+)1.69 in chrF++. It is surprising considering Medium has less than half of Single Ref’s n_source.
This suggests that when one has limited source and target texts, augmenting the number of references per source text can improve performance similarly to increasing the number of source text single target text pairs. However, it is unknown to what extent this observation will still hold when the references are synthetically generated with paraphrasers instead of high-quality human references in the literary domain.
6.2 Unfiltered All vs. Medium+High All
Although Unfiltered All has more training instances, its MT performance across different models and metrics is worse than Medium+High All and Medium+High 20% (Table 2).
We perform the paired bootstrap resampling to test the statistical significance of our finding (Koehn, 2004) and find that Med+High All outperforms Unfiltered All across different MT metrics and LLMs (p-value < 0.05). Specifically, Medium+High All outperforms Unfiltered All by 0.3-0.5 BLEU, 0.2-0.9 COMET, and 0.17-0.2 chrF++ (Medium+High 20% by 0.25-0.32 chrF++). This implies that when fine-tuning an MT model with multiple references, it is beneficial to compute the sim_p and filter out the ones with low semantic similarity to achieve the optimal MT performance.
Furthermore, we observe that Medium+Low 10% and All perform worse than Unfiltered All across different models and metrics (Table 2). Too much divergence among references of a source text may create confusion and hinder the performance.
@@ 6.3 Performance difference between models
We mentioned in Section 5.3 that LLaMA-2-7B is sensitive to the batch size when fine-tuning, and overall the fine-tuning result is less stable than mT5. In Table 2, we notice that BLEU, COMET, and chrF++ scores all seem less stable under LLaMA-2-7B, e.g. for Medium, increasing #instance resulted in a decrease in these metrics when #instance went from 57,987 to 76,938.
BLEU $\text{COMET}_{22}^{\text{DA}}$ chrF++
mT5 LLaMA-2-7B mT5 LLaMA-2-7B mT5 LLaMA-2-7B
medium + 1.0 31.7 28.9 72.0 71.3 44.59 42.46
medium + 0.9-1.0 31.7 27.8 71.9 70.3 44.55 40.75
medium + 0.85-0.9 31.7 31.0 71.9 72.9 44.52 44.70
Table 3: Ablation study on fine-tuning with the highly similar paraphrases (§6.5). 1.0 is pure repetition, 0.9-1.0 is the paraphrases with trivial differences, and 0.85-0.9 is non-trivial but non-meaningful paraphrases. All experiments have the same amount of source texts, #source = 57,987. For each MT metric, $\dagger$ denotes statistical significance for experiments using mT5, and $\ddagger$ denotes statistical significance for experiments using LLaMA-2-7B. We tested the statistical significance (p-value < 0.05) using the paired bootstrap resampling method (Koehn, 2004).
LLaMA-2’s MT ability is still not well understood: Xu et al. (2024) observe that for LLaMA-2-7B, increasing training data, under COMET, the MT performance declines, while under BLEU, it improves. From their observation, they suggested that for LLaMA-2-7B, increasing parallel data might dilute its pre-existing knowledge.
Although it is unknown what caused such unstable performance for LLaMA-2-7B, overall our observation in Section 6.2 still applies: using Medium+High is better than the unfiltered dataset, despite unfiltered dataset has larger #instance.

### v2

@@ Abstract
While a source sentence can be translated in many ways, most machine translation (MT) models are trained with only a single reference. Previous work has shown that using synthetic paraphrases can improve MT. This paper investigates best practices for employing multiple references by analyzing the semantic similarity among different English translations of world literature in the Par3 dataset (Thai et al., 2022). We classify the semantic similarity between paraphrases into three levels: low, medium, and high, and fine-tune three different models (mT5-large, LLaMA-2-7B, and Opus-MT) for literary MT tasks. Across different models, holding the total training instances constant, single-reference but more source texts only marginally outperforms multiple-reference with half of the source texts. Moreover, when fine-tuning an LLM, using paraphrases with medium and high semantic similarity outperforms an unfiltered dataset, with improvements in BLEU (0.3–0.5), COMET (0.1–0.9), and chrF++ (0.17–0.32). Our code is publicly available on GitHub. [^1] https://github.com/swsiwu/multi_ref_literary_MT
@@ 1 Introduction
In this paper, we investigate two questions: when holding the number of fine-tuning data constant, is using multiple references better than just one target reference? What kind of paraphrase should be incorporated to achieve the best performance?
To address these questions, we construct datasets that vary in the number of source texts, the number of references per source text (single vs. multiple references), and, for multiple-reference datasets, in the level of semantic similarity among the references. We use Paragram-SP (Wieting et al., 2022), an efficient and effective model for measuring semantic similarity. We categorize paraphrases into three groups based on their semantic similarity: low, medium, and high, where medium has the most interesting and meaningful variations among the references for a source text. We then fine-tune three different models: mT5-large, LLaMA-2-7B, and Opus-MT, using the aforementioned datasets on MT tasks and compare their performances against baseline and commercial models: No Language Left Behind (NLLB Team et al., 2022) and Google Translate.
@@
  • A dataset of source texts with medium and high semantic similarity among references can outperform an unfiltered dataset even when the unfiltered dataset has more training instances. (Section 6.2)
  • In contrast, training with source texts that have medium and low semantic similarity among references results in worse performance than an unfiltered dataset. (Section 6.2)
@@ 2 Related work
Khayrallah et al. (2020) use simulated paraphrases to improve MT performance on low-resource languages. They focus on the general effect of augmenting training data using synthetic paraphrases when training data is scarce. Our work further studies the semantic similarity of paraphrases on MT. It is worth noting that the languages present in our dataset are not low-resource languages, and instead of training a language model from scratch, we fine-tune three different types of models: mT5 (a text-to-text encoder-decoder model), LLaMA-2-7B (an autoregressive language model), and Opus-MT (a bidirectional transformer model for machine translation).
@@ 3 Measuring semantic similarity
Figure 2: The distribution of sim_p scores for the source texts in the Par3 corpus. Paraphrases with higher scores exhibit greater semantic and syntactic similarity. Upon manual inspection, we observe that the paraphrases within the red and blue vertical lines have meaningful variations: different enough to be interesting, yet similar enough to retain their nature as paraphrases.
@@
e.g., single vs double quotation marks, or word replacement Unfiltered
@@ 4 Datasets
We use #source to denote the number of unique source texts, and #instance denotes the total number of training instances, e.g., 3 source texts each has 3 reference translations, will be denoted as #source = 3, and #instance = 9.
We create four different types of datasets:
  • Single: Each source text has one reference.
  • Medium: Each source text has multiple references, where the semantic similarity between references is medium.
  • Unfiltered: Each source text has multiple references, with semantic similarity between references varying across the dataset.
  • Medium+Low, Medium+High: Each source text has multiple references. These datasets are constructed by starting with the entire Medium dataset and then incrementally adding 10% of Low or High similarity references until references are exhausted.
For each dataset type above except the last type, we create subsets with different numbers of sources or instances by performing random sampling within each dataset category.
The source texts and human references are from Par3 (Thai et al., 2022). [^2] Par3 dataset can be downloaded from https://github.com/katherinethai/par3 The source languages include Portuguese, Swedish, Italian, Czech, Japanese, Classical Chinese, Norwegian, German, Russian, and French. All references are only in English; In other words, the language pairs in this dataset are to-English language pairs. Details on the composition of the Medium dataset compared to the Unfiltered dataset are provided in Appendix 5.
@@ 5 Experiments
We investigate the impact of varying the number of source texts, references, and semantic similarity of reference translations on MT tasks in the literary domain. Using datasets described in Section 4, we fine-tune three different models for translation: mT5-large, LLaMA-2-7B, and Opus-MT.
@@ 5.1 Models
: An encoder-decoder model that covers 101 languages (Xue et al., 2021) including all languages in Par3. It has about 1.2 billion parameters. While not specifically trained for translation, it can be fine-tuned for multilingual translation and other downstream tasks. We fine-tuned this model using Hugging Face Trainer class. [^3] https://huggingface.co/google/mt5-large
@@
: An English-centric autoregressive transformer model with limited multilingual capabilities (Touvron et al., 2023). Studies have shown that LLaMA-2 can translate high-resource non-English language with some fine-tuning due to the presence of the corresponding language in the training dataset. LLaMA-2 is also able to adapt to a new low-resource language with some fine-tuning. All the languages in this paper are in the LLaMA-2 training data for >= 0.05%, except that the Chinese in Par3 is classical Chinese, which shares mostly the same characters as modern Chinese but has different grammar. It would be interesting to observe the translation performance after fine-tuning on our datasets given that non-English data is only a minuscule part of LLaMA-2 training data.
Opus-MT
: An open-source multilingual machine translation model. The specific version we use is opus-mt-mul-en, which supports multiple source languages and English as the only target language. All models that are part of the Opus-MT project (Tiedemann et al., 2023; Tiedemann and Thottingal, 2020) are trained on the OPUS corpus using the Marian framework (Junczys-Dowmunt et al., 2018), with an architecture similar to BART (Lewis et al., 2020). Unlike mT5-large and LLaMA-2-7B, Opus-MT is not a large language model, and it performs worse than the previous two. Similar to mT5, it’s also encoder-decoder architecture, but it leverages both bidirectional and autoregressive training objectives. This model is available on Hugging Face. [^4] https://huggingface.co/Helsinki-NLP/opus-mt-mul-en
@@ 5.2 Evaluation metrics
(Popović, 2017): An MT evaluation metric that is based on character and word-level n-gram recall and precision. We include this metric as Zouhar et al. (2024) find that COMET can struggle in domains outside WMT, e.g., literary domain.
#SRC #INST BLEU $\text{COMET}_{22}^{\text{DA}}$ chrF++
NLLB-1.3B 27.3 68.9 41.62
Google Translate 35.8 72.6 48.02
mT5 LLaMA Opus mT5 LLaMA Opus mT5 LLaMA Opus
single 50k 50k 29.9 30.1 23.2 71.5 70.3 66.6 43.44 42.08 38.28
medium 21k 50k 29.8 30.0 23.0 71.3 72.8 65.9 43.20 44.06 37.99
single 58k 58k 30.4 30.2 23.9 71.5 72.1 67.0 43.81 42.86 38.63
medium 24k 58k 29.8 30.2 23.1 71.3 72.4 66.0 43.36 43.79 38.06
single 77k 77k 31.3 29.4 24.3 71.8 71.3 67.1 44.33 42.81 38.97
medium 32k 77k 30.6 28.2 23.7 71.6 70.2 66.5 43.73 41.12 38.50
Table 2: To assess the efficiency of using multiple references, we keep the total number of training instances (#instance) constant. We compare the performance between each Single (where each source text has only one single human reference) and Medium (where each source text has more than two human references). A paired bootstrap resampling statistical significance test is performed on each pair, with the Single in each pair serving as the baseline (Koehn, 2004). All results are statistically significant ($p<0.01$). Across different models and metrics, Single is only marginally better than Medium, despite having at least twice as many source texts (#source).
#SRC #INST BLEU $\text{COMET}_{22}^{\text{DA}}$ chrF++
NLLB-1.3B 27.3 68.9 41.62
Google Translate 35.8 72.6 48.02
mT5 LLaMA Opus mT5 LLaMA Opus mT5 LLaMA Opus
unfiltered 77k 181k 31.9 30.8 25.3 71.9 72.1 67.5 44.67 44.11 39.54
medium 50k 118k 31.5 30.3 24.5 71.9 71.6 67.1 43.81 43.76 38.95
medium + low 58k 138k 31.1 29.4 24.5 71.7 71.6 67.0 44.16 41.29 38.97
medium + low 59k 140k 31.1 30.1 24.5 71.7 71.7 67.1 44.11 43.64 39.01
medium + high 58k 137k 32.0 29.2 24.7 72.0 70.3 67.4 44.77 42.39 39.21
medium + high 66k 156k 32.2 30.8 25.3 72.1 72.6 67.6 44.92 44.43 39.55
medium + high 68k 159k 32.4 31.1 25.1 72.1 73.0 67.5 44.84 44.31 39.47
Table 3: When working with a dataset where each source text has multiple references, using just Medium (where references have medium semantic similarity) can already yield results comparable to those of the unfiltered dataset. To assess the effect of adding low and high similarity references, we incrementally increase their proportion by 10% until exhausted. The results indicate that using all Medium + High references likely gives the best performance. We tested statistical significance using paired bootstrap resampling, with the unfiltered dataset as the baseline. All results are statistically significant ($p<0.01$) unless marked with $\dagger$.
@@ 5.3 Training details
Given many mT5 and LLaMA already have strong language abilities and only require fine-tuning for MT downstream tasks, we consider fine-tuning as a more efficient approach than training from scratch, especially since our dataset is English-centric (all reference translations are in English) and these LLMs usually perform better when translating from non-English language into English.
For fine-tuning, each training instance is a source text and one of its human translations, and the training and validation datasets are randolmized, e.g., (source text A, target human translation $A_{1}$), (source text B, target human translation $B_{3}$), (source text C, target human translation $C_{1}$), etc.
@@
For LLaMA-2-7B, we used LoRA-tuning (Hu et al., 2022) with Meta’s LLaMA-recipe. [^8] https://llama.meta.com/llama-downloads/, [^9] https://github.com/meta-llama/llama-recipes We structured each training instance as "source_text ###> target text" following the fine-tuning method from Zhang et al. (2023). Depending on the dataset size, training took 8 to 24 hours.
For Opus-MT, it already has translation ability. We fine-tuned the model using the Hugging Face Trainer class for it to adopt to the literary domain. For each experiment, we fine-tuned the model for 10 epochs with a batch size of 32 on a V100-SXM2 GPU. Depending on the dataset size, fine-tuning took between 1 and 3 hours.
For all models, we use BLEU on the validation set to select the best epoch. Unlike mT5, we find that LLaMA-2’s performance is significantly influenced by fine-tuning hyperparameters (±$\pm 5$ BLEU). Additionally, we observe that increasing the global batch size improves performance as the training data grows. We report the best performance achieved with three global batch sizes (8, 16, 32) for LLaMA-2-7B.
@@ 6.1 Single reference vs multiple references
This is not, to our knowledge, an idea that has been previously explored. In Table 2, we compare Medium vs Single under different total number of instances. When #instance is the same, Single on average only marginally outperforms Medium across different models and metrics: ranging 0.0–0.7 in BLEU, (-)2.5–(+)1.1 in COMET, and (-)1.98–(+)1.69 in chrF++. It is surprising considering Medium has less than half of Single’s #source. For each language model, we tested the statistical significance (p < 0.01) of the result by comparing the Medium and Single datasets with the same #instance, using Single as the baseline. The statistical significance was assessed using paired bootstrap resampling method (Koehn, 2004). [^10] For BLEU and chrF++: https://github.com/mjpost/sacrebleu,  [^11] For COMET: https://github.com/Unbabel/COMET
Our results suggest that, with limited source and target texts, increasing the number of references per source text can improve performance similarly to increasing the number of source-target pairs. However, it remains unclear whether this holds when references are synthetically generated by paraphrasers rather than high-quality human references in the literary domain.
6.2 Unfiltered vs. filtered datasets
Despite having more training instances, Unfiltered generally performs worse than Medium+High (#source $=66k,68k$) across models and metrics, with LLMs showing more notable improvements than the NMT model (Table 3).
We test the statistical significance of our result (Koehn, 2004) and find that Med+High (#source $=66k,68k$) outperforms Unfiltered across different MT metrics and models (p-value < 0.01). Specifically, with mT5 and LLaMA, Medium+High (#source $=68k$) outperforms Unfiltered by 0.3–0.5 BLEU, 0.2–0.9 COMET, and 0.17–0.2 chrF++ (Medium+High #source $=66k$ by 0.25–0.32 in chrF++). This implies that when fine-tuning an MT model with multiple references, it is beneficial to compute the sim_p and filter out the ones with low semantic similarity to achieve the optimal MT performance.
Furthermore, we observe that Medium+Low perform worse than Unfiltered across different models and metrics (Table 3). Too much divergence among references of a source text may create confusion and hinder the performance.
BLEU $\text{COMET}_{22}^{\text{DA}}$ chrF++
mT5 LLaMA Opus mT5 LLaMA Opus mT5 LLaMA Opus
medium + 1.0 31.7 28.9 24.6 72.0 71.3 67.3 44.59 42.46 39.16
medium + 0.9-1.0 31.7 27.8 24.8 71.9 70.3 67.3 44.55 40.75 39.29
medium + 0.85-0.9 31.7 31.0 24.9 71.9 72.9 67.4 44.52 44.70 39.30
Table 4: Ablation study on fine-tuning with the highly similar paraphrases (§6.5). 1.0 is pure repetition, 0.9-1.0 is the paraphrases with trivial differences, and 0.85–0.9 is non-trivial but non-meaningful paraphrases. All experiments have the same number of source texts, #source = 58k. All results are statistically significant ($p<0.01$) unless marked with $\dagger$.
@@ 6.3 Performance difference between models
The three models we fine-tune for literary MT task differ in architecture: mT5-large and LLaMA-2-7B are LLMs, while Opus-MT is a much smaller NMT model. As expected, Opus-MT’s performance is generally worse than the LLMs. More interestingly, under the same dataset type, mT5-large and Opus-MT performance improves when #instance increases, whereas LLaMA-2-7B is more unstable. For example, Medium’s BLEU, COMET, and chrF++ scores do not improve as we increase #instance under LLaMA-2-7B.
One potential reason for this discrepancy is that, unlike mT5-large and Opus-MT, which are both multilingual models, LLaMA-2-7B as a primary monolingual model. Its MT ability is still not well understood (Xu et al., 2024). On the other hand, as we mentioned in Section 5.3, LLaMA-2-7B is sensitive to the batch size using LoRA-tuning, which can also account for the less stable performance.